OpenStudy (kc_kennylau):

@Zupari Can you give me some exercises related to mole calculation and the like? I am especially weak in this topic of Chemistry :/

OpenStudy (anonymous):

Well do you need some serious problems or normal ones?

OpenStudy (kc_kennylau):

As you like, b/c I'm learning Chemistry by myself.

OpenStudy (kc_kennylau):

Like the previous question that you answered

OpenStudy (kc_kennylau):

I just want to be fluent in this topic

OpenStudy (kc_kennylau):

(excuse me for my grammar mistakes of any kind, coz my brain ain't functioning properly now)

OpenStudy (anonymous):

Try this mlar masses of C6H12O6 and water are 180 and 18 respectively.Glucose : Water is 2 : 3 according to their masses. find the mole fraction of glucose

OpenStudy (anonymous):

Actually I am not studing chemistry in English. So even I dont know some words

OpenStudy (kc_kennylau):

What language is your native language?

OpenStudy (anonymous):

You probably dont know

OpenStudy (kc_kennylau):

Say it :)

OpenStudy (anonymous):

Its Sinhala

OpenStudy (kc_kennylau):

180*2 : 18*3 = 10*2 : 1*3 = 20 : 3

OpenStudy (kc_kennylau):

you're right on that haha

OpenStudy (kc_kennylau):

Am I correct?

OpenStudy (anonymous):

I am asking mole fraction of glucose. It means no. of glucose moles/Total moles

OpenStudy (kc_kennylau):

Wait, something wrong...

OpenStudy (kc_kennylau):

20:23 correct? I don't feel like it....

OpenStudy (anonymous):

Its like this. the ratio is given according to their masses.We should always try not take masses as variables.So we should convert masses into moles

OpenStudy (kc_kennylau):

So 2/180 : (2/180 + 3/18) ?

OpenStudy (anonymous):

C6H12O6 : H2O = 2 : 3 according to their masses By converting to moles = 2/180 : 3/18

OpenStudy (anonymous):

n=m/M

OpenStudy (kc_kennylau):

So I'm correct?

OpenStudy (anonymous):

2/180 : 3/18 NOT 2/180 : (2/180 + 3/18)

OpenStudy (anonymous):

masses should be divided by molar masses

OpenStudy (anonymous):

In glucose its 2/180 and in water its 3/18

OpenStudy (kc_kennylau):

But you say total moles...

OpenStudy (anonymous):

Now we have found glucose and water moles So glucose mole fraction is 2/180/ (2/180 + 3/18)

OpenStudy (kc_kennylau):

Oh, I see

OpenStudy (anonymous):

answer is 0.0625

OpenStudy (kc_kennylau):

Thanks!

OpenStudy (anonymous):

Hey current time in here is 10.22 PM. so i should probably go. I'll give u some more questions tommorrow

OpenStudy (anonymous):

Is that okay?

OpenStudy (kc_kennylau):

Okay, it's 0:55 for me now so I'll have to sleep anyways, see you :D

OpenStudy (anonymous):

Bye